# Neutrino mixing parameters and masses from $\Delta(96)\rtimes H_{CP}$ in the tri-direct CP approach

### v1

@@ 2.1 The tri-direct approach
[…]
where the fermion fields above are adopted to two-component fermion notation. In Eq. (2.7), $L$ and $E^{c}$ represent left-handed lepton doublets and right-handed charged leptons, respectively. The flavons $\xi_{\text{atm}}$ and $\xi_{\text{sol}}$ are the Standard Model and flavor symmetry singlets, whereas the flavons $\phi_{l}$, $\phi_{\text{sol}}$ and $\phi_{\text{atm}}$ can be either Higgs fields or combinations of the electroweak Higgs doublet together with flavons.
@@
where $\langle\phi_{\text{atm}}\rangle$, $\langle\phi_{\text{sol}}\rangle$, $\langle\xi_{\text{atm}}\rangle$ and $\langle\xi_{\text{sol}}\rangle$ denote the vacuum alignments of flavons $\phi_{\text{sol}}$, $\phi_{\text{atm}}$, $\xi_{\text{atm}}$ and $\xi_{\text{sol}}$, respectively. The Dirac neutrino mass matrix $m_{D}$ is defined in the convention $(\nu_{L})^{T}m_{D}N^{c}$. The light neutrino mass matrix derived via the seesaw mechanism can be expressed as
@@
where $m_{l}$, the charged lepton mass matrix, is defined using the convention $(E^{c})^{T}m_{l}E_{L}$. The representation $\bm{r}$ corresponds to faithful, irreducible three-dimensional representations of the finite group $\Delta(96)$. With $G_{l}$ defined, the general structure of $m^{\dagger}_{l}m_{l}$ is directly formulated using Eq. (2.12). The unitary matrix $U_{l}$, satisfying $U^{\dagger}_{l}m^{\dagger}_{l}m_{l}U_{l}=\text{diag}(m^{2}_{e},m^{2}_{\mu},m^{2}_{\tau})$, governs the diagonalization of this hermitian product. As revealed by Eq. (2.12), $U_{l}$ is determined through the eigenvalue decomposition relation:
@@
In the neutrino sector, our analysis focuses exclusively on the NO case, with analogous results for the IO case derivable through similar procedures. Within the atmospheric and solar neutrino sectors, the vacuum alignment directions $\bm{v}_{\text{atm}}$ and $\bm{v}_{\text{sol}}$ of flavons $\phi_{\text{atm}}$ and $\phi_{\text{sol}}$ must remain invariant under their respective residual symmetry groups $G_{\text{atm}}\rtimes H^{\text{atm}}_{CP}$ and $G_{\text{sol}}\rtimes H^{\text{sol}}_{CP}$, respectively. This requirement translates to the constraints:
@@
constitutes a eigenvector with zero eigenvalue $m_{1}=0$ of the light neutrino mass matrix $m_{\nu}$ in Eq. (2.9). Hence, the normalized version $\bm{\hat{v}}_{\text{fix}}\equiv\bm{v}_{\text{fix}}/\sqrt{\bm{v}^{\dagger}_{\text{fix}}\bm{v}_{\text{fix}}}$ corresponds to the first column of the diagonalizing matrix $U_{\nu}$ satisfying $U^{T}_{\nu}m_{\nu}U_{\nu}=\text{diag}(0,m_{2},m_{3})$. As a consequence, the first column of the PMNS matrix is fixed to be $U^{\dagger}_{l}\bm{\hat{v}}_{\text{fix}}$.
@@
$U_{\nu 1}=\left(\bm{\hat{v}}_{\text{fix}},\bm{\hat{v}}^{*}_{\text{atm}},\bm{\hat{v}}^{\prime}_{\text{sol}}\right)\,,$ (2.16)
@@
Note that $U_{\nu 1}$ in Eq. (2.16) is non-unique and may include an additional unitary rotation in the $(23)$ plane in the case of $\bm{v}^{\dagger}_{\text{sol}}\bm{v}_{\text{atm}}\neq 0$. Consequently, the neutrino mass matrix $m_{\nu}$ can be reduced to a simpler structure:
$m_{\nu}^{\prime}=U_{\nu 1}^{T}m_{\nu}U_{\nu 1}=\begin{pmatrix}0&0&0\\ 0&y&z\\ 0&z&w\end{pmatrix}\,,$ (2.18)
@@
$U_{\nu 2}=\begin{pmatrix}1&0&0\\ 0&\cos\theta e^{i(\psi+\rho)/2}&\sin\theta e^{i(\psi+\sigma)/2}\\ 0&-\sin\theta e^{i(-\psi+\rho)/2}&\cos\theta e^{i(-\psi+\sigma)/2}\end{pmatrix}\,.$ (2.20)
The nonzero neutrino masses $m_{2}$ and $m_{3}$, along with the parameters $\theta$, $\psi$, $\rho$ and $\sigma$ are analytically determined by the parameters $y$, $z$ and $w$ in the simplified mass matrix $m_{\nu}^{\prime}$, as established in Refs. [15, 76, 77], and the explicit expressions of them can be found in appendix B.
In short, within the residual mixing pattern $\{G_{l},G_{\text{atm}}\rtimes H^{\text{atm}}_{CP},G_{\text{sol}}\rtimes H^{\text{sol}}_{CP}\}$, the explicit forms of $U_{l}$, $\bm{v}_{\text{atm}}$ and $\bm{v}_{\text{sol}}$ can be straightforwardly determined. However, due to the undetermined charged lepton mass hierarchy in the tri-direct CP approach, the PMNS matrix remains defined only up to independent row permutations. Consequently, the lepton mixing matrix for this residual mixing pattern is resolved modulo row reordering. Thus the general expression for the mixing matrix is derived as
@@
$\begin{array}[]{lll}P_{123}=\begin{pmatrix}1&~{}0~{}&0\\ 0&~{}1~{}&0\\ 0&~{}0~{}&1\end{pmatrix},{}{}&~{}~{}P_{132}=\begin{pmatrix}1&~{}0~{}&0\\ 0&~{}0~{}&1\\ 0&~{}1~{}&0\end{pmatrix},{}{}&~{}~{}P_{213}=\begin{pmatrix}0&~{}1~{}&0\\ 1&~{}0~{}&0\\ 0&~{}0~{}&1\end{pmatrix},\\ &&\\[-10.0pt] P_{231}=\begin{pmatrix}0&~{}1~{}&0\\ 0&~{}0~{}&1\\ 1&~{}0~{}&0\end{pmatrix},{}{}&~{}~{}P_{312}=\begin{pmatrix}0&~{}0~{}&1\\ 1&~{}0~{}&0\\ 0&~{}1~{}&0\end{pmatrix},{}{}&~{}~{}P_{321}=\begin{pmatrix}0&~{}0~{}&1\\ 0&~{}1~{}&0\\ 1&~{}0~{}&0\end{pmatrix}\,.\end{array}$ (2.22)
@@
where $c_{ij}\equiv\cos\theta_{ij}$, $s_{ij}\equiv\sin\theta_{ij}$, $\delta_{CP}$ is the Dirac CP violation phase and $\beta$ is the Majorana CP phase. For quantitative characterization of CP violation, two basis independent quantities are introduced: $J_{CP}$ [85] corresponding to the Dirac phase and $I_{1}$ [86, 87, 88, 89, 90] associated with the Majorana phase, defined respectively through:
@@
[…]
@@ 3 Lepton mixing predictions from $\Delta(96)\rtimes H_{CP}$
This section conducts a comprehensive model independent analysis of viable lepton mixing patterns within the framework of $\Delta(96)\rtimes H_{CP}$ by using the tri-direct CP approach. Through systematic scanning of all possible remnant symmetry combinations $\{G_{l},G_{\text{atm}}\rtimes H^{\text{atm}}_{CP},G_{\text{sol}}\rtimes H^{\text{sol}}_{CP}\}$, and the corresponding resulting predictions for lepton mixing parameters and neutrino masses are obtained. Unlike model dependent approaches requiring symmetry breaking mechanisms via charged scalar fields, our method focuses solely on residual symmetry structures. The PMNS matrix emerges from misalignment among the three residual symmetries, bypassing explicit vacuum alignment dynamics.
@@ 3.1 Lepton mixing patterns for $\phi_{\text{atm}},\phi_{\text{sol}}\sim\mathbf{\bar{3}_{0}}$
Observable $\text{bf}\pm 1\sigma$ $3\sigma$ region Observable $\text{bf}\pm 1\sigma$ $3\sigma$ region
$\sin^{2}\theta_{13}$ $0.02215^{+0.00056}_{-0.00058}$ $[0.02030,0.02388]$ $\frac{\Delta m^{2}_{21}}{10^{-5}\text{eV}^{2}}$ $7.49^{+0.19}_{-0.19}$ $[6.92,8.05]$
$\sin^{2}\theta_{12}$ $0.308^{+0.012}_{-0.011}$ $[0.275,0.345]$ $\frac{\Delta m^{2}_{31}}{10^{-3}\text{eV}^{2}}$ $2.513^{+0.021}_{-0.019}$ $[2.451,2.578]$
$\sin^{2}\theta_{23}$ $0.470^{+0.017}_{-0.013}$ $[0.435,0.585]$ $\frac{\Delta m^{2}_{21}}{\Delta m^{2}_{31}}$ $0.0298^{+0.00079}_{-0.00079}$ $[0.0268,0.0328]$
$\delta_{CP}/\pi$ $1.178^{+0.144}_{-0.228}$ $[0.689,2.022]$
Table 1: The global best fit values, $1\sigma$ ranges and $3\sigma$ ranges of mixing parameters and lepton mass ratios for NO neutrino mass spectrum. Here the experimental data and errors of the lepton mixing parameters and neutrino masses are obtained from NuFIT 6.0 with Super-Kamiokande atmospheric data [50].
@@
where the observables $O_{i}\pm\sigma_{i}$ represent the global best fit values and $1\sigma$ uncertainties for the six observable quantities including the three mixing angles $\sin^{2}\theta_{12}$, $\sin^{2}\theta_{13}$ and $\sin^{2}\theta_{23}$, the Dirac CP phase $\delta_{CP}$ and the mass squared differences $\Delta m^{2}_{21}$, $\Delta m^{2}_{31}$. In Eq. (3.5), $P_{i}$ represent the theoretical predictions for those six observable quantities which are formulated as functions of the input parameters $m_{a}$, $r$ and $\eta$. For each $(m_{a},r,\eta)$ parameter set, we compute both the predicted observables $P_{i}$ and the associated $\chi^{2}$ value. A comprehensive $\chi^{2}$ minimization analysis has been conducted across the parameter space.
After performing the $\chi^{2}$ analysis for all possible independent breaking patterns, we find only one phenomenologically interesting mixing patterns. The representative combination of residual flavor symmetries is $\{G_{l},G_{\text{atm}},G_{\text{sol}}\}=(Z_{3}^{ac},Z_{2}^{a^{2}b},Z_{3}^{ad^{2}})$, and both the atmospheric and solar neutrino sectors exhibit trivial residual CP transformations represented by the identity matrix $\mathbb{1}_{3}$. In this mixing pattern, the diagonalization matrix of charged lepton mass matrix $U_{l}$, the VEVs $\langle\phi_{\text{atm}}\rangle$ and $\langle\phi_{\text{sol}}\rangle$ are taken to be
@@
Obvious the solar and the reactor mixing angles satisfy the following sum rule
@@
which holds true for all TM1 mixing matrix. The lepton mixing matrix in Eq. (3.10) yields explicit expressions for the CP invariants:
@@
Under maximal atmospheric mixing $\theta_{23}=\pi/4$, the vanishing cotangent term $\cot 2\theta_{23}=0$ enforces $\cos\delta_{CP}=0$, thereby producing maximal CP violation with $\delta_{CP}=\pm\pi/2$. Figure 1 displays the contour plot of $\delta_{CP}/\pi$ within the $\sin^{2}\theta_{13}-\sin^{2}\theta_{23}$ parameter plane. The red regions correspond to viable parameter space where input variables $r$ and $\eta$ are sampled within permissible ranges, subject to experimental constraints ($3\sigma$) on mixing angles, Dirac CP phase and the ratio $\Delta m^{2}_{21}/\Delta m^{2}_{31}=m^{2}_{2}/m^{2}_{3}$. Remarkably, numerical simulations constrain $\delta_{CP}$ to a narrow interval near $1.45\pi$, demonstrating predictive power despite parametric freedom.
@@
where $m_{ee}$ refer to the neutrinoless double beta ($0\nu\beta\beta$) decay effective mass which is defined as
$m_{ee}=|m_{1}U^{2}_{PMNS,1}+m_{2}U^{2}_{PMNS,2}+m_{3}U^{2}_{PMNS,3}|\,.$ (3.17)
The predictions for mixing parameters in Eq. (3.16) exhibit excellent consistency with current experimental measurements, achieving a minimized $\chi^{2}$ function is $\chi^{2}_{\text{min}}=5.545$.
Furthermore, we conduct an extensive numerical study by treating the input parameters $m_{a}$, $r$ and $\eta$ as randomly sampled real numbers within the intervals $[0,1]\text{eV}$, $[0,100]$ and $[0,2\pi]$, respectively. For each parameter set, we compute the lepton mixing parameters and neutrino masses. The computed values of the six observable quantities in table 1 are constrained to lie in their $3\sigma$ confidence regions derived from recent global analyses of neutrino oscillation data [50]. Our analysis reveals the following experimentally compatible parameter ranges $m_{a}/\text{meV}\in[23.600,25.840]$, $r\in[0.338,0.381]$ and $\eta/\pi\in[0.571,0.643]$. The model predicts tightly constrained neutrino mass and mixing parameters:

### v2

@@ 2.1 The tri-direct approach
$-\frac{1}{2}x_{\text{atm}}\xi_{\text{atm}}N_{\text{atm}}^{c}N_{\text{atm}}^{c}-\frac{1}{2}x_{\text{sol}}\xi_{\text{sol}}N_{\text{sol}}^{c}N_{\text{sol}}^{c}+\text{h.c.}\,$,
where the fermion fields above are adopted to two-component fermion notation. In Eq. (2.7), $L$ and $E^{c}$ represent left-handed lepton doublets and right-handed charged leptons, respectively. The flavons $\xi_{\text{atm}}$ and $\xi_{\text{sol}}$ are the Standard Model and flavor symmetry singlets, whereas the flavons $\phi_{l}$, $\phi_{\text{sol}}$ and $\phi_{\text{atm}}$ can be either Higgs fields or combinations of the electroweak Higgs doublet together with flavons. Note that, in order to avoid cross term $N_{\text{sol}}^{c}N_{\text{atm}}^{c}$ in Eq. (2.7), the right-handed neutrinos $N_{\text{sol}}^{c}$ and $N_{\text{atm}}^{c}$ must be assigned to different singlets under $\Delta(96)$. Otherwise, the model would require an additional $Z_{n}$ ($n\geq 2$) symmetry, under which the two fields carry distinct charges.
@@
where $\langle\phi_{\text{atm}}\rangle$, $\langle\phi_{\text{sol}}\rangle$, $\langle\xi_{\text{atm}}\rangle$ and $\langle\xi_{\text{sol}}\rangle$ denote the vacuum alignments of flavons $\phi_{\text{atm}}$, $\phi_{\text{sol}}$, $\xi_{\text{atm}}$ and $\xi_{\text{sol}}$, respectively. The Dirac neutrino mass matrix $m_{D}$ is defined by the convention $(\nu_{L})^{T}m_{D}N^{c}$. The light neutrino mass matrix derived via the seesaw mechanism can be expressed as
@@
where $m_{l}$, the charged lepton mass matrix, is defined using the convention $(E^{c})^{T}m_{l}E_{L}$. The representation $\bm{r}$ corresponds to faithful, irreducible three-dimensional representation of the finite group $\Delta(96)$. With $G_{l}$ defined, the general structure of $m^{\dagger}_{l}m_{l}$ is directly formulated using Eq. (2.12). The unitary matrix $U_{l}$, satisfying $U^{\dagger}_{l}m^{\dagger}_{l}m_{l}U_{l}=\text{diag}(m^{2}_{e},m^{2}_{\mu},m^{2}_{\tau})$, governs the diagonalization of this hermitian product. As revealed by Eq. (2.12), $U_{l}$ is determined through the eigenvalue decomposition relation:
@@
In the atmospheric and solar neutrino sectors, the vacuum alignment directions $\bm{v}_{\text{atm}}$ and $\bm{v}_{\text{sol}}$ of flavons $\phi_{\text{atm}}$ and $\phi_{\text{sol}}$ must remain invariant under their respective residual symmetry groups $G_{\text{atm}}\rtimes H^{\text{atm}}_{CP}$ and $G_{\text{sol}}\rtimes H^{\text{sol}}_{CP}$, respectively. This requirement translates to the constraints:
@@
corresponds to the zero eigenvalue $m_{1}$=0 of the light neutrino mass matrix $m_{\nu}$ in the NO case, and to $m_{3}$=0 in the IO case. Hence, the normalized version $\bm{\hat{v}}_{\text{fix}}\equiv\bm{v}_{\text{fix}}/\sqrt{\bm{v}^{\dagger}_{\text{fix}}\bm{v}_{\text{fix}}}$ corresponds to the first column of the diagonalizing matrix $U_{\nu}$ satisfying $U^{T}_{\nu}m_{\nu}U_{\nu}=\text{diag}(0,m_{2},m_{3})$ for the NO case and the third column of the diagonalizing matrix $U_{\nu}$ satisfying $U^{T}_{\nu}m_{\nu}U_{\nu}=\text{diag}(m_{1},m_{2},0)$ for the IO case. As a consequence, the first (third) column of the PMNS matrix is fixed to be $U^{\dagger}_{l}\bm{\hat{v}}_{\text{fix}}$ for the NO (IO) case.
@@
$U_{\nu 1}=\left\{\begin{array}[]{lll}{(\bm{\hat{v}}_{\mathrm{fix}},\bm{\hat{v}}_{\mathrm{atm}}^{*},\bm{\hat{v}}_{\mathrm{sol}}^{\prime})}&{\mathrm{for}}&{\mathrm{NO}\,,}\\ {(\bm{\hat{v}}_{\mathrm{atm}}^{*},\bm{\hat{v}}_{\mathrm{sol}}^{\prime},\bm{\hat{v}}_{\mathrm{fix}})}&{\mathrm{for}}&{\mathrm{IO}\,.}\end{array}\right$. (2.16)
@@
Note that $U_{\nu 1}$ in Eq. (2.16) is non-unique: when $\bm{v}^{\dagger}_{\text{sol}}\bm{v}_{\text{atm}}\neq 0$, for the NO case, it may include a (23) rotation, while for the IO case, it may include a (12) rotation. Consequently, the neutrino mass matrix $m_{\nu}$ can be reduced to a simpler structure:
$m_{\nu}^{\prime}=U_{\nu 1}^{T}m_{\nu}U_{\nu 1}=\begin{cases}\begin{pmatrix}0&0&0\\ 0&y&z\\ 0&z&w\end{pmatrix}&\mathrm{for}\quad\mathrm{NO\,,}\\ \begin{pmatrix}y&z&0\\ z&w&0\\ 0&0&0\end{pmatrix}&\mathrm{for}\quad\mathrm{IO\,.}\end{cases}$ (2.18)
@@
$U_{\nu 2}=\left\{\begin{array}[]{ccc}\begin{pmatrix}1&0&0\\ 0&\cos\theta e^{i(\psi+\rho)/2}&\sin\theta e^{i(\psi+\sigma)/2}\\ 0&-\sin\theta e^{i(-\psi+\rho)/2}&\cos\theta e^{i(-\psi+\sigma)/2}\end{pmatrix}&\mathrm{for}&\mathrm{NO\,,}\\ \\ \begin{pmatrix}\cos\theta e^{i(\psi+\rho)/2}&\sin\theta e^{i(\psi+\sigma)/2}&0\\ -\sin\theta e^{i(-\psi+\rho)/2}&\cos\theta e^{i(-\psi+\sigma)/2}&0\\ 0&0&1\end{pmatrix}&\mathrm{for}&\mathrm{IO\,.}\end{array}\right$. (2.20)
For the NO (IO) case, the nonzero neutrino masses $m_{2}$ and $m_{3}$ ($m_{1}$ and $m_{2}$), along with the parameters $\theta$, $\psi$, $\rho$ and $\sigma$ are analytically determined by the parameters $y$, $z$ and $w$ in the simplified mass matrix $m_{\nu}^{\prime}$, as established in Refs. [15, 76, 77], and the explicit expressions of them can be found in appendix B.
In short, within the residual mixing pattern $\{G_{l},G_{\text{atm}}\rtimes H^{\text{atm}}_{CP},G_{\text{sol}}\rtimes H^{\text{sol}}_{CP}\}$, the explicit forms of $U_{l}$, $\bm{v}_{\text{atm}}$ and $\bm{v}_{\text{sol}}$ can be straightforwardly determined. However, due to the undetermined charged lepton mass hierarchy in the tri-direct CP approach, the PMNS matrix remains defined only up to independent row permutations. Consequently, the lepton mixing matrix for this residual symmetry pattern is resolved modulo row reordering. Thus the general expression for the mixing matrix is derived as
@@
[…]
@@
where $c_{ij}\equiv\cos\theta_{ij}$, $s_{ij}\equiv\sin\theta_{ij}$, $\delta_{CP}$ is the Dirac CP violation phase and $\beta$ is the Majorana CP phase. For quantitative characterization of CP violation, two basis independent quantities are introduced: $J_{CP}$ [86] corresponding to the Dirac phase and $I_{1}$ [87, 88, 89, 90, 91] associated with the Majorana phase, defined respectively by:
@@
$I_{1}$ $=$ $\left\{\begin{array}[]{lll}\Im{(U^{2}_{12}U^{*\,2}_{13})}=\frac{1}{4}\sin^{2}\theta_{12}\sin^{2}2\theta_{13}\sin(\beta+2\delta_{CP})&\text{for }&\text{NO}\,,\\ \Im{(U^{2}_{12}U^{*\,2}_{11})}=\frac{1}{4}\cos^{4}\theta_{13}\sin^{2}2\theta_{12}\sin\beta&\text{for}&\text{ IO}\,.\end{array}\right$. (2.26)
@@ 3 Lepton mixing predictions from $\Delta(96)\rtimes H_{CP}$
This section conducts a comprehensive model independent analysis of viable lepton mixing patterns within the framework of $\Delta(96)\rtimes H_{CP}$ by using the tri-direct CP approach. Through systematic scanning of all possible remnant symmetry combinations $\{G_{l},G_{\text{atm}}\rtimes H^{\text{atm}}_{CP},G_{\text{sol}}\rtimes H^{\text{sol}}_{CP}\}$, the corresponding resulting predictions for lepton mixing parameters and neutrino masses are obtained. Unlike model dependent approaches requiring symmetry breaking mechanisms via charged scalar fields, our method focuses solely on residual symmetry structures. The PMNS matrix emerges from misalignment among the three residual symmetries, bypassing explicit vacuum alignment dynamics.
@@
Note that if both $\phi_{\text{atm}}$ and $\phi_{\text{sol}}$ are assigned to the $\bm{\bar{3}_{0}}$ representation, then the right-handed neutrinos $N_{\text{atm}}^{c}$ and $N_{\text{sol}}^{c}$ must both transform as $\bm{1_{0}}$ under $\Delta(96)$. This assignment, however, necessitates introducing an additional $Z_{n}$ ($n\geq 2$) symmetry to forbid the unwanted cross term $N_{\text{atm}}^{c}N_{\text{sol}}^{c}$. The compensating advantage of this setup is that it requires only three input parameters, eliminating the need for the additional parameter $x$. When $\phi_{\text{atm}}$ and $\phi_{\text{sol}}$ are assigned to $\bm{\bar{3}_{0}}$ and $\bm{\bar{3}_{1}}$ respectively, the right-handed neutrinos $N_{\text{atm}}^{c}$ and $N_{\text{sol}}^{c}$ must correspondingly transform as $\bm{1_{0}}$ and $\bm{1_{1}}$. In this case, the cross term vanishes naturally, eliminating the need for an additional symmetry to distinguish the two fields. The main drawback, however, is the requirement of an additional real parameter $x$.
@@ 3.1 Lepton mixing patterns for $\phi_{\text{atm}},\phi_{\text{sol}}\sim\mathbf{\bar{3}_{0}}$
Observables NO IO
$\text{bf}\pm 1\sigma$ $3\sigma$ region $\text{bf}\pm 1\sigma$ $3\sigma$ region
$\sin^{2}\theta_{13}$ $02215^{+0.00056}_{-0.00058} [0.0$2030,0.02388][0.02030,0.02388] $02231^{+0.00056}_{-0.00056} [0.0$2060,0.02409][0.02060,0.02409]
$\sin^{2}\theta_{12}$ $0.308^{+0.012}_{-0.011}$ $[0.275,0.345]$ $0.308^{+0.012}_{-0.011}$ $[0.275,0.345]$
$\sin^{2}\theta_{23}$ $0.470^{+0.017}_{-0.013}$ $[0.435,0.585]$ $0.550^{+0.012}_{-0.015}$ $[0.440,0.584]$
$\delta_{CP}/\pi$ $1.178^{+0.144}_{-0.228}$ $[0.689,2.022]$ $522^{+0.122}_{-0.139} [1.1$17,1.861][1.117,1.861]
$\frac{\Delta m^{2}_{21}}{10^{-5}\text{eV}^{2}}$ $7.49^{+0.19}_{-0.19}$ $[6.92,8.05]$ $7.49^{+0.19}_{-0.19}$ $[6.92,8.05]$
$\frac{\Delta m^{2}_{3\ell}}{10^{-3}\text{eV}^{2}}$ $2.513^{+0.021}_{-0.019}$ $[2.451,2.578]$ $2.484^{+0.020}_{-0.020}$ [2.547,2.421][-2.547,-2.421]
$\Delta m^{2}_{21}/\Delta m^{2}_{3\ell}$ $0298^{+0.00079}_{-0.00079} [0.0$268,0.0328][0.0268,0.0328] $0302^{+0.00080}_{-0.00080} [0.0$333,0.0272][-0.0333,-0.0272]
Table 1: The global best fit values, $1\sigma$ ranges and $3\sigma$ ranges of mixing parameters and lepton mass ratios. Here the experimental data and errors of the lepton mixing parameters and neutrino masses are obtained from NuFIT 6.0 with Super-Kamiokande atmospheric data [50]. Note that the notation $\Delta m^{2}_{3\ell}$ denotes $\Delta m^{2}_{31}>0$ for the NO and $\Delta m^{2}_{32}<0$ for the IO.
@@
where the observables $O_{i}\pm\sigma_{i}$ represent the global best fit values and $1\sigma$ uncertainties for the six observable quantities including the three mixing angles $\sin^{2}\theta_{12}$, $\sin^{2}\theta_{13}$ and $\sin^{2}\theta_{23}$, the Dirac CP phase $\delta_{CP}$ and the mass squared differences $\Delta m^{2}_{21}$, $\Delta m^{2}_{31}$ ($\Delta m^{2}_{32}$) for the NO (IO) case. In Eq. (3.5), $P_{i}$ represents the theoretical predictions for those six observable quantities which are formulated as functions of the input parameters $m_{a}$, $r$ and $\eta$. For each $(m_{a},r,\eta)$ parameter set, we compute both the predicted observables $P_{i}$ and the associated $\chi^{2}$ value. A comprehensive $\chi^{2}$ minimization analysis has been conducted across the parameter space.
After performing the $\chi^{2}$ analysis for all possible independent breaking patterns, we find only one phenomenologically interesting mixing pattern. The representative combination of residual flavor symmetries is $\{G_{l},G_{\text{atm}},G_{\text{sol}}\}=(Z_{3}^{ac},Z_{2}^{a^{2}b},Z_{3}^{ad^{2}})$, and both the atmospheric and solar neutrino sectors exhibit trivial residual CP transformations represented by the identity matrix $\mathbb{1}_{3}$. In this mixing pattern, the diagonalization matrix of charged lepton mass matrix $U_{l}$, the VEVs $\langle\phi_{\text{atm}}\rangle$ and $\langle\phi_{\text{sol}}\rangle$ are taken to be
@@
Obviously, the solar and the reactor mixing angles satisfy the following sum rule
@@
which holds true for all TM1 mixing matrices. The lepton mixing matrix in Eq. (3.10) yields explicit expressions for the CP invariants:
@@
Under maximal atmospheric mixing $\theta_{23}=\pi/4$, the vanishing cotangent term $\cot 2\theta_{23}=0$ enforces $\cos\delta_{CP}=0$, thereby producing maximal CP violation with $\delta_{CP}=\pm\pi/2$. Figure 1 displays the contour plot of $\delta_{CP}/\pi$ within the $\sin^{2}\theta_{13}-\sin^{2}\theta_{23}$ parameter plane. The red region corresponds to viable parameter space where input variables $r$ and $\eta$ are sampled within permissible ranges, subject to experimental constraints ($3\sigma$) on mixing angles, Dirac CP phase and the ratio $\Delta m^{2}_{21}/\Delta m^{2}_{31}=m^{2}_{2}/m^{2}_{3}$. Remarkably, numerical simulations constrain $\delta_{CP}$ to a narrow interval near $1.45\pi$, demonstrating predictive power despite parametric freedom.
@@
where $m_{ee}$ refers to the neutrinoless double beta ($0\nu\beta\beta$) decay effective mass which is defined as
$m_{ee}=|m_{1}U^{2}_{PMNS,11}+m_{2}U^{2}_{PMNS,12}+m_{3}U^{2}_{PMNS,13}|\,.$ (3.17)
The predictions for mixing parameters in Eq. (3.16) exhibit excellent consistency with current experimental measurements, with a minimum $\chi^{2}$ value of $\chi^{2}_{\text{min}}=5.545$.
Given the high flavor symmetry breaking scale, it is necessary in principle to account for renormalization group (RG) running down to the electroweak scale. Comprehensive treatments of the RG evolution of lepton mixing parameters and non-zero neutrino masses can be found in Refs. [92, 93, 94]. These studies show that the corrections to the mixing parameters due to RG running are negligible, as they are significantly smaller than current experimental uncertainties. This behavior arises because, in the case of a massless lightest neutrino, the running lacks a strong enhancement mechanism. Therefore, our low-energy predictions remain robust against RG running effects.
Furthermore, we conduct an extensive numerical study by treating the input parameters $m_{a}$, $r$ and $\eta$ as randomly sampled real numbers within the intervals $0,1]\text{eV}$, $[0,100]$ and $0,2\pi]$, respectively. For each parameter set, we compute the lepton mixing parameters and neutrino masses. The computed values of the six observable quantities in table 1 are constrained to lie in their $3\sigma$ confidence regions derived from recent global analyses of neutrino oscillation data [50]. Our analysis reveals the following experimentally compatible parameter ranges $m_{a}/\text{meV}\in[23.600,25.840]$, $r\in[0.338,0.381]$ and $\eta/\pi\in[0.571,0.643]$. The model predicts tightly constrained neutrino mass and mixing parameters: